How would you find the remaining area of a circle with a rectangle in it measure 4 ft by 3 ft with all four corners touching the outer edge of the circle?
Answer by edjones(8007)

The diagonal of the rectangle is 5 because it is the hypotenuse of a 3-4-5 right triangle. The diagonal is also the diameter of the circle.
A=pi*r^2
=pi*2.5^2
=pi*6.25 ft^2
.
Ed